Beryllium sulfate is a hydrated compound whose formula can be written BeSO4 • xH2O, where z is the number of moles of H2O per mole of BeSO4. When a 3.284 g sample of this hydrate was heated to 130°C for 10 minutes, all of the water of hydration was lost, leaving 1.948 g of anhydrous beryllium sulfate. Calculate the value of x, and write the formula and name of the hydrate.


1
Expert's answer
2021-05-04T14:16:24-0400

Δm = 3.284 – 1.948 = 1.336 g

M(H2O) = 18 g/mol

n(H2O) "= \frac{1.336}{18} = 0.07422 \; mol"

M(BeSO4) = 105.07 g/mol

n(BeSO4) "= \frac{1.948}{105.07} = 0.01854 \;mol"

0.01854 : 0.07422 = 1: 4

x = 4

BeSO4 • 4H2O Beryllium sulfate tetrahydrate